zarkam21 (zarkam21):

Solve the system of equations. Enter your answer as an ordered pair. 3x-4y=-30 3x+2y=6

OpenStudy (anonymous):

elimination might be easier

zarkam21 (zarkam21):

I know I don't understand it

OpenStudy (anonymous):

3x - 4y = -30 --->(-1) 3x + 2y = 6 ------------ -3x + 4y = 30 (result of multiplying by -1) 3x + 2y = 6 -----------add 6y = 36 y = 36/6 y = 6 now sub 6 back in for y in either of the original equations to find x

zarkam21 (zarkam21):

Umm Does X= -2

zarkam21 (zarkam21):

@kelliegirl33

OpenStudy (kidrah69):

\[3x+2(6)=6\]